OpenStudy (dls):

The velocity of a particle is at any time related to the distance traveled by the relation V(x)=ax+b,where a is positive and b<=a/2.Which of the following is true for the motion?

OpenStudy (dls):

@Yahoo!

OpenStudy (dls):

The displacement of the particle at time t is x=b/a(e^at -1) The particle will experience retardation if b<0 The particle will be at rest at time t=0 Constant acceleration

OpenStudy (dls):

@ujjwal

OpenStudy (anonymous):

its not D for sure :P

OpenStudy (dls):

IDK the answer!

OpenStudy (anonymous):

oh wait.. i m wrong.. never mind whatever i said!

OpenStudy (dls):

take ur time :o

OpenStudy (anonymous):

ok the fourth answer is wrong cause if acceleration is zero.. then dv(x)/dt = zero which means ad(x)/dt = zero.. which implies velocity is zero!

OpenStudy (anonymous):

oh i think its this way the given velocity can be converted to a function of time v(t) = ax(t) + b hence dx(t)/dt = ax(t) + b.. then solve this differential equation!

OpenStudy (anonymous):

did you get it DLS?

OpenStudy (dls):

no.. :|

OpenStudy (anonymous):

I got the answer..!! but you need to solve that differential equation !!.. i just converted the function to time domain!!! you know the differential forms of accerelation and velocity right?

OpenStudy (dls):

let me try..

OpenStudy (anonymous):

ok.. tell me what you get the solution of the D.E

OpenStudy (dls):

a?

OpenStudy (anonymous):

No .. you have to integrate it!!

OpenStudy (dls):

:o

OpenStudy (dls):

ax^2+b/2 thats what i did before

OpenStudy (anonymous):

no no!!... in the equation in time domain.. you have to put x(t) on one side .. so rearrange the equation as dx(t)/(ax(t)+b) = d(t) then integrate.. both sides.. you get log(ax+b)/a = t or ax+b = e^at or x = (e^at-b)/a

OpenStudy (dls):

WAIT IDK ALL THAT !

OpenStudy (anonymous):

ok sorry .. .. so no calculus?? no differentiation and integration??!

OpenStudy (dls):

only diff and integration i know but that domain rerranginng and log stuff IDK

OpenStudy (anonymous):

thats how you solve a differential equation.. and all i did was put it interms of time :P.. velocity was given interms of displacement.. so i put displacement in terms of time :P

OpenStudy (dls):

please use equation editor for better explanation!

OpenStudy (anonymous):

sorry.. first tell me.. do you know how to solve a differential equation?

OpenStudy (dls):

example?

OpenStudy (anonymous):

example if i say \[d(x)/dt = xt\] find x ?

OpenStudy (dls):

\[\frac{dt}{dy}=2t^{4} +3t^{2} = 8t^{3}+6t\] I can solve this :p

OpenStudy (dls):

how do we solve that?

OpenStudy (anonymous):

ok its solved this way.. first rearrange the terms \[d(x)/dt=xt\] bring x terms one side and t terms on one side so \[d(x)/x = t dt\] now integrate \[\int\limits_{}^{}d(x)/x = \int\limits_{}^{}tdt\]